OpenStudy (anonymous):

Chase wants to factor x2 + 10x + 25 by grouping; however, Paige says it is a special product and can factor a different way. Using complete sentences, explain and demonstrate how both methods will result in the same factors.

OpenStudy (anonymous):

I keep forgetting how to split the middle term.

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

To get quadratics of the form x^2+ax+b in the form (x+c)(x+d) we need to find a c and d such that: c*d=b c+d=a In this case we need to find c*d=25 c+d=10 Can you find a combination of c and d which satisfy those equations?

OpenStudy (superhelp101):

this is a perfect square trinomial so it would factor like (x+5)(x+5) because what numbers multiply to get 25 and add to get 10. That would x+5

OpenStudy (anonymous):

Do I look at factors of 25 or 10?

OpenStudy (superhelp101):

that one was a special products (x+5)^2

OpenStudy (superhelp101):

factors of 25 are 5 and 5 so the numbers 5 and 5 multiply to get 25 and they also add to get 10. That is what we are looking for

OpenStudy (anonymous):

So how do I do the factor by grouping way

OpenStudy (superhelp101):

um that's the only way I was taught @tom982 do you think you know

OpenStudy (anonymous):

After a quick Google I know what you mean by a 'grouping way' now. \[x^2+10x+25 = (x^2+5x) + (5x+25) = x(x+5)+5(x+5) = (x+5)(x+5)\]

OpenStudy (anonymous):

Thank you so much. I don't know why I keep forgetting how to split the middle term.

OpenStudy (anonymous):

No problemo, happy to help!